### v2

@@ 3 Preliminaries
In this section, we begin by introducing the notations used in this paper. Building upon this foundation, we then elaborate on the threat model and problem formulation related to the specific attack scenario of the SIA. For preliminary analysis, we investigate a heuristic technique known as GraphCopy to assess the viability of the proposed SIA scenario and establish a foundational benchmark.
@@ 4 Methodology
In this section, we investigate an optimization-based approach for generating subgraphs to realize the SIA scenario. While extensive research has focused on targeting either node classification or link prediction models independently [13], the exploitation of the interaction between these two distinct model types to achieve an adversarial objective remains largely unexplored. Moreover, achieving dual adversarial goals within the SIA scenario presents significant technical challenges. Intuitively, generating links is more likely when there is a close resemblance in the node representations between the target victim nodes and the nodes of the aggressor subgraph, whereas a greater disparity in these representations increases the chances of the victim nodes being misclassified. As a result, a more deliberate subgraph optimization technique is essential to balance two conflicting objectives and identify the optimal subgraph for SIA. To this end, we introduce the LiSA framework, as depicted in Fig. 2.
@@ 5.1 Experiment Settings
5.1.1 Datasets
To evaluate the effectiveness of our attack methods, we perform experiments on five distinct datasets. These datasets include Cora and PubMed, both citation networks, Amazon-Photo and Computers, co-purchase networks from a popular e-commerce platform, and FacebookPagePage, representing a social network graph.
5.1.2 Experimental Verification
We conduct an experimental verification of our LiSA framework using GAE and VGAE [8] for link recommender, along with GCN [9], SGC [15], and GraphSAGE [7] as node classifiers. Our attack methodology is applied across all datasets with the presumption that the target user would establish links with nodes of top three linking scores in the entire poisoned graph. To assess the feasibility of our methods, we randomly select 1000 nodes from each original graph as target victim nodes. We then generate subgraphs and execute attacks on these nodes individually. This approach allows us to compute both the link success rate (LSR) and the overall attack success rate (ASR). The LSR is defined as the proportion where at least one adversarial edge is generated between the target node and subgraph, while the ASR refers to the misclassfication rate on these victim nodes. For the link recommender, we use 85% of the edges for training, with half serving as positive edge labels with negative sampling for both subgraph optimization and attack verification. For node classification, 5% of the nodes are randomly chosen for training labels. These labels are utilized in subgraph optimization and the pretraining of a node classifier to identify the second highest output class as the attack label. To maintain a consistent basis for comparison across different datasets, we employ a subgraph comprising five nodes for each target victim node attack. However, we tailor hyperparameters such as the number of subgraph edges $n_{E}$, and the balance parameters $\alpha$ and $\beta$ for each dataset, ensuring optimal performance and adaptability to the specific characteristics of each dataset. We also compare the LiSA framework with baseline of GraphCopy in the SIA scenario, and an additional baseline of conventional NIA scenario. This particular version of NIA incorporates the probability of successful link creation, aiming to demonstrate the influence of link generation on the success of NIAs.
@@ 5.2 Attack Results
5.3 Impacts of Subgraph Size and User Added Links
We conduct a series of experiments on the Amazon-Photo dataset to investigate how subgraph size impact the attack. Newly generated links are chosen based on top linking scores in the experiments. Our initial observation reveal a decline in LSR as the subgraph size increased, a trend depicted in Fig. 4(a). This decline likely stems from the challenges associated with optimizing larger subgraphs, especially when adjusting both structure and features. However, we noted a stabilization in LSR when the subgraph size exceeds 5. Another interesting finding related to the response of ASR to the number of links added by users, as illustrated in Fig. 4(b). With the addition of a single link, we observe that the ASR is highest when the subgraph consist of a single node, decreasing as the subgraph size expands, possibly due to the corresponding reduction in LSR. On the other hand, with five links added, the ASR improve as the subgraph grow to five nodes, before experiencing a slight decrease with further increases in subgraph size. This phenomenon can be attributed to the rise in aggressor nodes linked to the target, enhancing effectiveness of the attack. However, exceeding the subgraph size with available links prompts the link recommender to seek top-scoring nodes from the original graph for link formation, which strengthens the resilience of target node to the attack [3]. Taking into account these effects, we found that the optimal ASR of 10 links was attained when the subgraph size was 7.
(a) LSR (%)
(b) ASR (%)
5.4 Subgraph Analysis
(a) Cora
(b) PubMed
Figure 5: Subgraph Analysis.
@@ Appendix 0.D NIA Settings, Results and Analysis
Appendix 0.E Parameter Sensitivity Analysis
@@
Appendix 0.F Subgraph Training Time Analysis

### v3

@@ 4 Methodology
In this section, we investigate an optimization-based approach for generating subgraphs. In the SIA scenario, achieving dual adversarial goals within the SIA scenario presents significant technical challenges. Intuitively, generating links is more likely when there is a close resemblance in the node representations between the target victim nodes and the nodes of the aggressor subgraph, whereas a greater disparity in these representations increases the chances of the victim nodes being misclassified. As a result, a more deliberate subgraph optimization technique is essential to balance two conflicting objectives and identify the optimal subgraph for SIA. To this end, we introduce the LiSA framework, as depicted in Fig. 2.
@@ 5.1 Experiment Settings
To evaluate the effectiveness of our attack methods, we perform experiments on five distinct datasets. These datasets include Cora and PubMed, both citation networks, Amazon-Photo and Computers, co-purchase networks from a popular e-commerce platform, and FacebookPagePage, representing a social network graph. We conduct an experimental verification of our LiSA framework using GAE and VGAE [7] for link recommender, along with GCN [8], SGC [13], and GraphSAGE [6] as node classifiers. Our attack methodology is applied across all datasets with the presumption that the target user would establish links with nodes of top three linking scores in the entire poisoned graph. To assess the feasibility of our methods, we randomly select 1000 nodes from each original graph as target victim nodes. We then generate subgraphs and execute attacks on these nodes individually. This approach allows us to compute both the link success rate (LSR) and the overall attack success rate (ASR). The LSR is defined as the proportion where at least one adversarial edge is generated between the target node and subgraph, while the ASR refers to the misclassfication rate on these victim nodes. For the link recommender, we use 85% of the edges for training, with half serving as positive edge labels with negative sampling for both subgraph optimization and attack verification. For node classification, 5% of the nodes are randomly chosen for training labels. These labels are utilized in subgraph optimization and the pretraining of a node classifier to identify the second highest output class as the attack label. To maintain a consistent basis for comparison across different datasets, we employ a subgraph comprising five nodes for each target victim node attack. However, we tailor hyperparameters such as the number of subgraph edges $n_{E}$, and the balance parameters $\alpha$ and $\beta$ for each dataset, ensuring optimal performance and adaptability to the specific characteristics of each dataset. We also compare the LiSA framework with baseline of GraphCopy in the SIA scenario, and an additional baseline of conventional NIA scenario. This particular version of NIA incorporates the probability of successful link creation, aiming to demonstrate the influence of link generation on the success of NIAs.
@@ Appendix 0.D NIA Settings, Results and Analysis
Appendix 0.E Impacts of Subgraph Size and User Added Links
We conduct a series of experiments on the Amazon-Photo dataset to investigate how subgraph size impact the attack. Newly generated links are chosen based on top linking scores in the experiments. Our initial observation reveal a decline in LSR as the subgraph size increased, a trend depicted in Fig. 5(a). This decline likely stems from the challenges associated with optimizing larger subgraphs, especially when adjusting both structure and features. However, we noted a stabilization in LSR when the subgraph size exceeds 5. Another interesting finding related to the response of ASR to the number of links added by users, as illustrated in Fig. 5(b). With the addition of a single link, we observe that the ASR is highest when the subgraph consist of a single node, decreasing as the subgraph size expands, possibly due to the corresponding reduction in LSR. On the other hand, with five links added, the ASR improve as the subgraph grow to five nodes, before experiencing a slight decrease with further increases in subgraph size. This phenomenon can be attributed to the rise in aggressor nodes linked to the target, enhancing effectiveness of the attack. However, exceeding the subgraph size with available links prompts the link recommender to seek top-scoring nodes from the original graph for link formation, which strengthens the resilience of target node to the attack. Taking into account these effects, we found that the optimal ASR of 10 links was attained when the subgraph size was 7.
(a) LSR (%)
(b) ASR (%)
Appendix 0.F Subgraph Analysis
(a) Cora
(b) PubMed
Figure 6: Subgraph Analysis.
Appendix 0.G Parameter Sensitivity Analysis
@@
Appendix 0.H Subgraph Training Time Analysis